OpenStudy (anonymous):

Solve: f^-1(3) when f(x)=2x+3/5 Can these be explained/work shown for me please?

OpenStudy (anonymous):

3 = 2x+3/5--solve for x, so subtract 3/5 to both sides 3-3/5=2x---combine like terms 12/5 = 2x --- divide both sides by 2 6/5 = x

OpenStudy (anonymous):

So the final answer is 1.2?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Alright, thank you for explaining. I've been stuck on this for abit because of the negative power.

OpenStudy (anonymous):

f^-1(x) mean the inverse of the function

OpenStudy (anonymous):

Thank you very much ^^. Medal awarded and fanned.

OpenStudy (anonymous):

f(x) = 2x+3/5 is the original function f^-1(x) is asking you for the inverse of the function so if there are no given values for x, you need to find the inverse function but if there is a value for x, the easiest and fastest was to solve it what i did at the top

OpenStudy (anonymous):

No Problem :)